\begin{thebibliography}{99}

\bibitem{Bal88}
K.~Ball.
\newblock Logarithmically conave functions and sections of convex sets in $\mathbb{R}^n$.
\newblock {\em Studia Math.} \textbf{88} (1988), 69--84.

\bibitem{BNT16}
K.~Ball, P.~Nayar, T.~Tkocz.
\newblock A reverse entropy power inequality for log-concave random vectors.
\newblock {\em Studia Math.} \textbf{235} (2016), 17--30.

\bibitem{BM12}
S. Bobkov, M. Madiman. 
\newblock Reverse Brunn-Minkowski and reverse entropy power inequalities for convex measures. 
\newblock {\em J. Funct. Anal.} \textbf{262} (2012), 3309-3339.

% \bibitem{BLL76}
% H. J. Brascamp, E. H. Lieb, J. M. Luttinger. 
% \newblock A general rearrangement inequality for multiple integrals. 
% \newblock {\em J. Funct. Anal.} \textbf{17} (1974), 227-237.


\bibitem{Bus49}
H.~Busemann.
\newblock A theorem on convex bodies of the Brunn-Minkowski type.
\newblock {\em Proc. Nat. Acad. Sci. USA} \textbf{35} (1949), 27--31.

\bibitem{Cho74}
K. M. Chong. 
\newblock Some extensions of a theorem of Hardy, Littlewood and P\'olya and their applications,
\newblock {\em Canad. J. Math.} \textbf{26} (1974), 1321-1340.

\bibitem{CZ94}
T.~M. Cover, Z.~Zhang.
\newblock On the maximum entropy of the sum of two dependent random variables.
\newblock {\em IEEE Trans. Inform. Theory} \textbf{40} (1994), 1244--1246.

\bibitem{Gar02}
R.~J. Gardner.
\newblock The Brunn-Minkowski inequality
\newblock {\em Bull. Amer. Math. Soc.} \textbf{39} (2002), 355-405

\bibitem{GG99}
R.~J. Gardner, A.~A. Giannopoulos.
\newblock $p$-cross-section bodies.
\newblock {\em Indiana Univ. Math. J.} \textbf{48} (1999), 593--613.

\bibitem{HLP29}
G. H. Hardy, J. E. Littlewood, G. P\'olya. 
\newblock Some simple inequalities satisfied by convex functions.
\newblock {\em Mess. of Math.} \textbf{58} (1929), 145--152.

\bibitem{HSTV22}
H. Huang, B. Slomka, T. Tkocz, B. Vritsiou. 
\newblock Improved bounds for Hadwiger's covering problem via thin-shell estimates.
\newblock {\em J. Eur. Math. Soc.} (JEMS) \textbf{24} (2022), 1431--1448.

\bibitem{Li18}
J. Li. 
\newblock R\'enyi entropy power inequality and a reverse, 
\newblock {\em Studia Math.} \textbf{242} (2018), 303-319.

\bibitem{MMX17}
M.~Madiman, J.~Melbourne, P.~Xu.
\newblock Forward and reverse entropy power inequalities in convex geometry.
\newblock In: Carlen, E., Madiman, M., Werner, E. (eds) Convexity and Concentration. The IMA Volumes in Mathematics and its Applications, vol 161. Springer, New York, NY. 

\bibitem{Mil86}
V.~D. Milman.
\newblock In{\'e}galit{\'e} de brunn-minkowski inverse et applicationsa la
  th{\'e}orie locale des espaces norm{\'e}s.
\newblock {\em C. R. Acad. Sci. Paris S\'er. I Math.} \textbf{302} (1986), 25--28.

\bibitem{MPC23}
J. Melbourne, G. Palafox-Castillo.
\newblock A discrete complement of Lyapunov's inequality and its information theoretic consequences.
\newblock {\em Ann. Appl. Probab.}  \textbf{33} (2023), 4322-4340.

\bibitem{MT21}
J. Melbourne, T. Tkocz.
\newblock Reversal of R\'enyi entropy inequalities under log-concavity.
\newblock {\em IEEE Trans. Inform. Theory} \textbf{67} (2021), 45-51.


\bibitem{RS57}
C. A. Rogers, G. C. Shephard.
\newblock The difference body of a convex body. 
\newblock {\em Arch. Math} \textbf{8} (1957), 220–233.

\bibitem{Rog87}
B. A. Rogozin. 
\newblock An estimate for the maximum of the convolution of bounded densities.
\newblock {\em Teor. Veroyatnost. i Primenen.} \textbf{32} (1987), 53–61.

\bibitem{Sha48}
C. E. Shannon.
\newblock A mathematical theory of communication.
\newblock {\em Bell System Tech. J.} \textbf{27} (1948), 379--423, 623--656.


\bibitem{Sim11}
B. Simon.
\newblock {\em Convexity: An Analytic Viewpoint}.
\newblock Cambridge Tracts in Mathematics, vol. 187. Cambridge: Cambridge University Press, 2011.


\bibitem{Sta59}
A.~J. Stam.
\newblock Some inequalities satisfied by the quantities of information of Fisher and Shannon.
\newblock {\em Inform. and Control} \textbf{2} (1959), 101--112.

\bibitem{WM14}
L. Wang, M. Madiman.
\newblock Beyond the Entropy Power Inequality, via Rearrangements.
\newblock {\em IEEE Trans. Inf. Theory} \textbf{60} (2014), 5116--5137.


\end{thebibliography}

\documentclass[12pt, reqno]{amsart}
\usepackage{mathrsfs}
\usepackage{amsfonts,amsmath,dsfont,amssymb,amsthm,stmaryrd,bbm,color,comment}
\usepackage[hidelinks]{hyperref}
\usepackage{appendix}
\usepackage[pdftex]{graphicx}
\usepackage{bbm}
\usepackage{array}
\newcolumntype{L}{>{\centering\arraybackslash}m{2.5cm}}
\usepackage{multirow}
\usepackage{xcolor}
\usepackage[percent]{overpic}
\usepackage{algorithm}
\usepackage{algpseudocode}
\usepackage{mathtools}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{conj}[thm]{Conjecture}

\theoremstyle{remark}
\newtheorem{rmk}[thm]{Remark}
\numberwithin{equation}{section}

\newcommand{\supp}{\mathrm{supp}}

\newcommand{\R}{\mathbb{R}}
\newcommand{\N}{\mathbb{N}}
\newcommand{\Z}{\mathbb{Z}}

\def\P{{\mathbb P}}

\begin{document}
\setcounter{page}{1}

\color{black}{

\centerline{}
\centerline{}
\title[A reverse EPI for i.i.d. log-concave random variables]{A reverse entropy power inequality for i.i.d. log-concave random variables}
\author[Zhen Fu, and Jiange Li]{Zhen Fu, and Jiange Li} 
\address{(Z.F, J.L) Institute for Advanced Study in Mathematics, Harbin Institute of Technology, China}
\email{zhenfu@stu.hit.edu.cn, jiange.li@hit.edu.cn}


\begin{abstract}
We show that $h_\infty(X+Y)\leq h_\infty(Z+W)$, where $X, Y$ are independent log-concave random variables, and $Z, W$ are exponential random variables having the same respective $\infty$-R\'enyi entropies. Analogs for integer-valued monotone log-concave random variables are also obtained. Our main tools are decreasing rearrangement, majorization, and the change of measure.
\end{abstract} 

\maketitle


\section{Introduction}
Let $X$ be a random vector taking values in $\R^d$. Suppose that it has density $f$ with respect to the Lebesgue measure. The classical \textit{differential entropy} (also called \textit{Boltzmann-Shannon entropy}) of $X$ is defined as
$$
h(X) =-\int_{\R^d} f(x)\log f(x)dx.
$$
The celebrated \textit{entropy power inequality} (or EPI for short) of Shannon \cite{Sha48} (and Stam \cite{Sta59}) states that for independent random vectors $X$ and $Y$ in $\R^d$ such that the entropies of $X, Y$, and $X+Y$ exist, it holds that
\begin{equation}\label{eq:epi}
e^{\frac{2}{d}h(X+Y)}\geq e^{\frac{2}{d}h(X)}+e^{\frac{2}{d}h(Y)},
\end{equation}
where equality holds if and only if $X$ and $Y$ are Gaussian random vectors with proportional covariance matrices. The EPI found its original applications in the study of channel capacity in information theory. It has been known very well that EPI is closely related to geometric and functional inequalities in a variety of mathematical fields, maybe most notably the Brunn-Minkowski inequality (or BMI for short) in convex geometry. We refer the interested reader to \cite{Gar02, MMX17} for profound connections among EPI, BMI, and many other geometric and functional inequalities.

Convexity (or concavity) underpins the validity of the reversal of many geometric and functional inequalities. A function $f:\R^d\to \R_+$ is called \textit{log-concave} if it can be written as $f(x)=e^{-V(x)}$, where $V: \R^d\to \R\cup\{\infty\}$ is a convex function. We say that a random vector $X$ taking values in $\R^d$ is log-concave provided that its density is a log-concave function. (We also say that $X$ has log-concave distribution). The class of log-concave functions can be thought of as functional analogs of convex bodies. They are prominent objects studied in convex geometry, high dimensional probability and statistics, as well as theoretical computer science. The class of log-concave distributions include a wide range of important probability distributions, such as the Gaussian distributions, uniform distributions on convex bodies, and exponential distributions. 

The reversal of EPI was first addressed by Bobkov and Madiman \cite{BM12}. They showed that for independent log-concave random vectors, under linear volume preserving maps, inequality \eqref{eq:epi} can be reversed at the cost of a constant factor on the r.h.s of the inequality. This can be thought of as a functional analog of Milman's reverse BMI for convex bodies \cite{Mil86}. It is necessary to place log-concave random vectors ``at right positions" via linear volume preserving maps for the reversed EPI to hold. We are interested in the sharp reverse EPI for \textit{independent and identically distributed} (or i.i.d. for short) log-concave random vectors. In this case, linear volume preserving maps are not needed and the question can be phrased as follows.

\begin{conj} [Folklore] \label{conj:folklore}
Let $X$ and $Y$ be i.i.d. log-concave random vectors taking values in $\R^d$. The entropy increment $h(X+Y)-h(X)$ is maximized when $X$ and $Y$ have exponential distributions. 
\end{conj}

Clearly, the entropy increment is affine invariant. An earlier result of Cover and Zhang \cite{CZ94} shows that log-concavity and identical distribution (without independence between $X$ and $Y$) yield the upper bound $d\log 2$ of the entropy increment. Huang, Slomka, Tkocz, and Vritsiou (\cite{HSTV22}, Section 6) observed that the improvement of Cover and Zhang's bound for Conjecture \ref{conj:folklore} will lead to important progress on Hadwiger's covering problem. Ball, Nayar, and Tkocz \cite{BNT16} studied the entropy increment for a two dimensional log-concave random vector $(X, Y)$, whose marginals have identical differential entropy (but not necessarily independent). This variant of reverse EPI and its generalizations are entropic analogs of Busemann's convexity theorem of intersection bodies \cite{Bus49} and the convexity of $p$-cross-section bodies conjectured by Gardner and Giannopoulos \cite{GG99} (the end of Section 5, and the discussion in \cite{Li18}).

Aforementioned forward and reverse EPIs have been studied for general R\'enyi entropies. For $p\in (0, 1)\cup (1, \infty)$, the order $p$ R\'enyi entropy of $X$ is defined as 
\begin{equation}\label{eq:renyi}
h_p(X)=\frac{1}{1-p}\log \int_{\R^d}f(x)^pdx.
\end{equation}
For $p\in\{0, 1, \infty\}$, by taking limits, we have 
$h_0(X)=\log |\supp\{f\}|$, $h_\infty(X) =-\log \|f\|_\infty$, and $h_1(X) =h(X)=-\int_{\R^d} f(x)\log f(x)dx$. Here, we denote by $|\supp\{f\}|$ the Lebesgue measure of the support of $f$, and $\|f\|_\infty$ is the essential supremum of $f$, and $h_1(X)$ is the classical differential entropy $h(X)$ of $X$. 

It is suspected that a transition will occur in the increment of R\'enyi entropy. More precisely, \textit{there exists an absolute constant $p_0\in (0, 1)$ such that $h_p(X+Y)-h_p(X)$ is maximized by exponential random vectors for $p\geq p_0$, and is maximized by uniform random vectors taking values in the cube $[0, 1]^d$ for $0<p\leq p_0$}. It is perhaps worth to point out that the increment of R\'enyi entropy of order $p>1$ (respectively, $0<p<1$) boils down to the minimization (respectively, maximization) of $\|f\ast f\|_p/\|f\|_p$ over the class of integrable log-concave functions. This can be thought of as a special type of reverse Young's convolution inequality.  Our main result reads as follows.

\begin{thm}\label{thm:reverse-rogozin}
Let $X$ and $Y$ be independent real-valued log-concave random variables. Let $Z$ and $W$ be independent exponential random variables such that $h_\infty(Z)=h_\infty(X)$ and $h_\infty(W)=h_\infty(Y)$. We have
\begin{align*}
h_\infty(X+Y)\leq h_\infty(Z+W).
\end{align*}
A discrete analog holds for integer-valued monotone log-concave random variables.
\end{thm} 

\begin{rmk}
If $X_1, \cdots, X_n$ are independent real-valued random variables, and $U_1, \cdots, U_n$ are independent uniform random variables such that $h_\infty(U_i)=h_\infty(X_i)$ for $i=1,\cdots, n$, Rogozin \cite{Rog87} proved that
$$
h_\infty(X_1+\cdots+X_n)\geq h_\infty(U_1+\cdots+U_n).
$$
Hence, our result can also be seen as a reverse Rogozin type result for two log-concave random variables. However, we are not able to establish analogs of Theorem \ref{thm:reverse-rogozin} for three or more log-concave random variables or in multi-dimensions.
\end{rmk}

Another related problem is the characterization of log-concave distributions that maximize the R\'enyi entropic increment $h_p(X-Y)-h_p(X)$. In this case, it is suspected that exponential distributions are always maximizers. This resembles the Rogers-Shephard inequality in convex geometry \cite{RS57}. In a certain sense, the symmetry of $X-Y$ (implicitly) makes this problem more tractable than its counterpart of the $X+Y$ case. Melbourne and Tkocz (\cite{MT21}, Theorem VI.1) showed that the increment of R\'enyi entropy of order $p\geq 2$ is indeed maximized by exponential random vectors. A discrete analog was obtained by Melbourne and Palafox-Castillo  (\cite{MPC23}, Theorem 2.11) for integer-valued monotone log-concave random variables.


\section{Rearrangement and majorization}

Rearrangement and majorization are classical tools for establishing sharp geometric and functional inequalities. This section provides the necessary preliminaries on decreasing rearrangement and majorization theory in one dimension. Readers already familiar with rearrangement theory are encouraged to skip this section.

\subsection{Decreasing rearrangement}
Let $A\subseteq \R$ be a Borel measurable set with Lebesgue measure $|A|$. The \textit{decreasing rearrangement} of $A$ is defined as $A^\downarrow =[0, |A|)$. The following properties of decreasing rearrangement may be scattered in the literature. We provide proof of them for the convenience of the reader.

\begin{prop}\label{prop:rearrange propt-set}
Let $A, B\subseteq\R$ be Borel measurable sets.
\begin{enumerate}
\item For all $\lambda\geq 0$, it holds that $(\lambda A)^\downarrow=\lambda A^\downarrow$.
\item  It holds that $A^\downarrow+B^\downarrow \subseteq (A+B)^\downarrow$.
\item  Let $\{A_n\}_{n=1}^\infty$ be a sequence of measurable sets such that $A_n\subseteq A_{n+1}$ for all $n=1, 2\cdots$. It holds that
$(\cup_{n=1}^\infty A_n)^\downarrow=\cup_{n=1}^\infty A_n^\downarrow$.
\end{enumerate}
\end{prop}

\begin{proof} 
Statement (1) readily follows from the definition. By definition, $A^\downarrow=[0, |A|)$, $B^\downarrow=[0, |B|)$, and $(A+B)^\downarrow =[0, |A+B|)$. The Brunn-Minkowski inequality in one dimension yields $|A+B|\geq |A|+|B|$. Hence, $A^\downarrow+B^\downarrow=[0,|A|+|B|) \subseteq (A+B)^\downarrow$. This proves statement (2).
 
By definition, $A_n^\downarrow=[0, |A_n|)$. Since $A_n\subseteq A_{n+1}$, we have
$$
\cup_{n=1}^\infty A_n^\downarrow=\big[0, \lim_{n\rightarrow\infty} |A_n|\big)=\left[0, |\cup_{n=1}^\infty A_n|\right)=(\cup_{n=1}^\infty A_n)^\downarrow.
$$
This proves the last statement.
\end{proof}

Let $f: \R\rightarrow \R\cup\{\pm \infty\}$ be a Borel measurable function such that all super level sets $\{f>\lambda\}$ have finite measures (which is called \textit{vanishing at infinity}). We define the \textit{decreasing rearrangement} $f^\downarrow: \R_+\rightarrow \R\cup\{\pm \infty\}$ as 
\begin{equation}\label{eq:mr-f}
f^\downarrow(x)=\sup\left\{\lambda\in\R: x\in \{f>\lambda\}^\downarrow\right\}.
\end{equation}

\begin{prop}\label{prop:rearrange propt-function}
Let $f: \R\rightarrow\R\cup\{\pm \infty\}$ be a Borel measurable function.
\begin{enumerate}
\item For all $\lambda\in\R$, it holds that $\{f^\downarrow>\lambda\}=\{f>\lambda\}^\downarrow$. (Hence, this gives an equivalent definition of the decreasing rearrangement $f^\downarrow$).
\item If $f$ is concave, then $f^\downarrow$ is also concave. Consequently,  if $f$ is log-concave, then $f^\downarrow$ is also log-concave.
\end{enumerate}
\end{prop}

\begin{proof}
For each $x\in\{f^\downarrow>\lambda\}$, by the definition of $f^\downarrow$ in \eqref{eq:mr-f}, there exists $\lambda_x>\lambda$ such that $x\in \{f>\lambda_x\}^\downarrow\subseteq \{f>\lambda\}^\downarrow$. This shows that $\{f^\downarrow>\lambda\}\subseteq \{f>\lambda\}^\downarrow$. To see the other direction, we apply Proposition \ref{prop:rearrange propt-set} (statement (3)) to write
$$
 \{f>\lambda\}^\downarrow=\left(\cup_{n=1}^\infty \{f>\lambda+1/n\}\right)^\downarrow=\cup_{n=1}^\infty \{f>\lambda+1/n\}^\downarrow.
$$
Hence, for each $x\in \{f>\lambda\}^\downarrow$, there exists $n\in\N$ such that $x\in \{f>\lambda+1/n\}^\downarrow$. By the definition of $f^\downarrow$ in \eqref{eq:mr-f}, this yields $f^\downarrow(x)\geq \lambda+1/n>\lambda$, and hence, $x\in \{f^\downarrow>\lambda\}$. This completes the proof of the first statement.

Suppose $f$ is concave. This is equivalent to the following set-theoretic inequality.  For any $\lambda_0, \lambda_1\in \R$ and $t\in [0, 1]$, it holds that
\begin{equation}\label{eq:concave}
(1-t)\{f>\lambda_0\}+t\{f>\lambda_1\}\subseteq \{f >(1-t)\lambda_0+t\lambda_1\}.
\end{equation}
Next we prove the inclusion inequality \eqref{eq:concave} for $f^\downarrow$. This will yield the concavity of $f^\downarrow$. We apply statement (1) of the current proposition and Proposition \ref{prop:rearrange propt-set} (statements (1) and (2))  to obtain
\begin{align*}
(1-t)\{f^\downarrow>\lambda_0\}+t\{f^\downarrow>\lambda_1\} &= (1-t)\{f>\lambda_0\}^\downarrow+t\{f>\lambda_1\}^\downarrow\\
&= ((1-t)\{f>\lambda_0\})^\downarrow+(t\{f>\lambda_1\})^\downarrow\\
&\subseteq ((1-t)\{f>\lambda_0\}+t\{f>\lambda_1\})^\downarrow\\
& \subseteq \{f>(1-t)\lambda_0+t\lambda_1\}^\downarrow\\
&=\{f^\downarrow >(1-t)\lambda_0+t\lambda_1\}.
\end{align*}

Suppose $f(x)=e^{V(x)}$ is log-concave, i.e., $V(x)$ is concave. Write $f^\downarrow(x)=e^{\tilde{V}(x)}$. For all $\lambda\in\R$, we have
$$
\{\tilde{V}>\lambda\}=\{f^\downarrow>e^{\lambda}\}=\{f>e^{\lambda}\}^\downarrow=\{V>\lambda\}^\downarrow.
$$
Using the first statement, we conclude that $\tilde{V}=V^\downarrow$. Since $V$ is concave, $V^\downarrow$, and hence, $\tilde{V}$ are concave. (One can show the concavity of $\tilde{V}$ by establishing the corresponding set-theoretic inequality). This yields the log-concavity of $f^\downarrow$.
\end{proof}

The following is a reverse Hardy-Littlewood type inequality. 
\begin{lem}\label{lem:rearrangement}
Let $f, g$ be  non-negative functions supported on $[a, b]$. Then it holds
$$
\int_a^b f(x)g(x)dx\geq \int_0^{b-a} f^\downarrow(x)g^\downarrow(b-a-x)dx.
$$
\end{lem}

\begin{proof}
Write $f(x)=\int_0^{f(x)}d\lambda$, $g(x)=\int_0^{g(x)}d\sigma$ and apply Fubini-Tonelli's theorem to obtain
\begin{align*}
\int_a^b f(x)g(x)dx &= \int_0^\infty\int_0^\infty\int_a^b\mathbbm{1}_{\{f>\lambda\}}(x)\mathbbm{1}_{\{g>\sigma\}}(x)dxd\lambda d\sigma\\
&=\int_0^\infty\int_0^\infty|\{f>\lambda\}\cap \{g>\sigma\}|d\lambda d\sigma.
\end{align*}
Similarly, we have
\begin{align*}
\int_0^{b-a} f^\downarrow(x)g^\downarrow(b-a-x)dx &=\int_0^\infty\int_0^\infty|\{f^\downarrow>\lambda\}\cap (b-a-\{g^\downarrow>\sigma\})|d\lambda d\sigma\\
&=\int_0^\infty\int_0^\infty|\{f>\lambda\}^\downarrow\cap (b-a-\{g>\sigma\}^\downarrow)|d\lambda d\sigma,
\end{align*}
where the second equality follows from  Proposition \ref{prop:rearrange propt-function} (statement (1)). Then the statement can be proved by applying the following inequality to $A=\{f>\lambda\}, B=\{g>\sigma\}$. For any measurable sets $A, B\subseteq [a, b]$, it holds that
$$
|A^\downarrow \cap (b-a-B^\downarrow)|\leq |A\cap B|.
$$
We assume that $A^\downarrow \cap (b-a-B^\downarrow)\neq \emptyset$. Note $A^\downarrow=[0, |A|)$ and $b-a-B^\downarrow=[b-a-|B|, b-a)$. Hence, we have
\begin{align*}
|A^\downarrow \cap (b-a-B^\downarrow)| &=|A|-(b-a-|B|)\\
&=|A|+|B|-(b-a)\\
&\leq |A|+|B|-|A\cup B|\\
&=|A\cap B|.
\end{align*}
This completes the proof.
\end{proof}


\subsection{Majorization}

Let $f, g: \R\to \R_+$ be two integrable functions such that $$\int_{\R}f(x)dx=\int_{\R}g(x)dx\textcolor{red}{.}$$ We say that $f$ is \textit{majorized} by $g$ if it holds for all $t\geq 0$ that
\begin{equation}\label{eq:maj-def}
\int_{\R}(f(x)-t)_+dx\leq \int_{\R}(g(x)-t)_+dx.
\end{equation}
For $t\geq 0$, we write $m_f(t)=|\{x\in\R: f(x)> t\}|$ and define $m_g(t)$ in a similar manner. Then we can rewrite \eqref{eq:maj-def} as
$$
\int_t^\infty m_f(s)ds\leq \int_t^\infty m_g(s)ds.
$$ 

A real-valued function $f$ on $\R$ is called \textit{unimodal} if there exists a point $x_0$ such that $f$ is non-decreasing for $x<x_0$ and non-increasing for $x>x_0$. Clearly, log-concave functions are unimodal. The following result provides a sufficient condition for the majorization of $g$ over $f$. 

\begin{prop}\label{prop:maj-suf}
Suppose that for any Borel measurable set $A$ of finite measure, there exists a Borel measurable set $B$ with equal measure as $A$ such that
\begin{equation}\label{eq:maj-suf}
\int_Af(x)dx\leq \int_Bg(x)dx.
\end{equation}
Then $f$ is majorized by $g$. If $f$ and $g$ are unimodal, it suffices to verify the above inequality for intervals.
\end{prop}

\begin{proof}
For any given $t\in\R_+$, we write $A=\{f>t\}$. Then there exists $B\subseteq\R$ of the same measure as $A$ such that
\begin{align*}
\int_{\R} (f(x)-t)_+dx &=\int_A (f(x)-t)dx\\
&\leq  \int_B (g(x)-t)dx\\
& \leq\int_B (g(x)-t)_+dx\\
&\leq \int_{\R} (g(x)-t)_+dx.
\end{align*}
This shows the majorization of $g$ over $f$. 

Note that the integral of a unimodal function over all measurable sets of fixed measure attains its maximum for an interval. Therefore, if $f$ is unimodal, it is sufficient to evaluate the integral on the left-hand side of inequality \eqref{eq:maj-suf} over intervals. Moreover, when $g$ is also unimodal, one often considers the integral on the right-hand side of inequality \eqref{eq:maj-suf} over intervals, even though considering general Borel measurable sets with the same measure as  $A$ would suffice.
\end{proof}

he following characterization of the majorization relation is well known. A preliminary version was first proved by Hardy, Littlewood, and P\'olya \cite{HLP29} and various extensions were discussed in Chong \cite{Cho74}. For the proof, we refer the reader to Theorem 2.1 in \cite{Cho74} and Theorem 15.27 in \cite{Sim11}; a related result can also be found in Proposition 7.3 of \cite{WM14}.

\begin{prop}\label{prop:major}
Let $f$ and $g$ be non-negative functions on $\R$. Then $f$ is majorized by $g$ if and only if for all non-negative increasing convex functions $\phi: \R_+\to \R_+$ with $\phi(0)=0$ it holds that
\begin{equation}\label{eq:maj}
\int_{\R}\phi(f(x))dx\leq \int_{\R}\phi(g(x))dx.
\end{equation}
\end{prop}

\section{Main Results}

\begin{prop}\label{prop:2-ent-jump}
Let $X$ and $Y$ be i.i.d. log-concave random vectors in $\R^d$. We have
\begin{equation}\label{eq:2-ent-jump}
h_2(X+Y)\leq h_2(X)+d\log 2.
\end{equation}
Moreover, the inequality can not be improved.
\end{prop}

\begin{rmk}
One can check that equality of \eqref{eq:2-ent-jump} can be achieved by exponential and Laplace distributions with densities $\prod_{i=1}^d\lambda_ie^{-\lambda_ix_i}\mathbbm{1}_{\R_+}(x_i)$, and $\prod_{i=1}^d\frac{\lambda_i}{2}e^{-\lambda_i|x_i|}$, respectively,  where $\lambda_i>0$ for $i=1, \cdots, d$.
\end{rmk}

\begin{proof}
Let $\tilde{X}$ and $\tilde{Y}$ be independent copies of $X$ and $Y$, respectively. Let $f$ be the density of $X-\tilde{X}$. Since $f$ is symmetric and log-concave, we have
$$
\|f\ast f\|_\infty=(f\ast f)(0)=\int_{\R^d}f(x)^2dx\geq f(0)\int_{\R^d}f(2x)dx=\frac{f(0)}{2^d}.
$$ 
This can be rewritten as
$$
h_\infty((X-\tilde{X})+(Y-\tilde{Y}))\leq  h_\infty(X-\tilde{X})+d\log 2.
$$
For any i.i.d. log-concave random vectors $U$ and $V$, it holds that $h_\infty(U-V)=h_2(U)$. This observation, together with the inequality above, implies that 
\begin{align}\label{eq:symmetrization}
h_2(X+Y) &=h_\infty((X+Y)-(\tilde{X}+\tilde{Y})) \notag\\
&=h_\infty((X-\tilde{X})+(Y-\tilde{Y})) \\
&\leq h_\infty(X-\tilde{X})+d\log 2 \notag\\
&= h_2(X)+d \log 2 \notag.
\end{align}
This completes the proof.
\end{proof}

\begin{rmk}\label{rmk:sigma2}
Note that $X-\tilde{X}$ and $Y-\tilde{Y}$ are symmetric and independent. Then, we can proceed with \eqref{eq:symmetrization} to obtain
\begin{align*}
h_2(X+Y) &=h_\infty((X-\tilde{X})+(Y-\tilde{Y}))\\
&=h_\infty((X-\tilde{X})-(Y-\tilde{Y}))\\
&=h_\infty((X-Y)-(\tilde{X}-\tilde{Y}))\\
&=h_2(X-Y).
\end{align*}
Hence, inequality \eqref{eq:2-ent-jump} is the same as the $p=2$ case of Theorem VI.1 of Melbourne and Tkocz \cite{MT21}. For any log-concave random vector $(X, Y)\in \R^2$, it was proved in \cite{Li18} that
$$
e^{h_2(X+Y)}\leq e^{h_2(X)}+e^{h_2(Y)},
$$
which implies the one-dimensional case of \eqref{eq:2-ent-jump} as long as $X, Y$ are jointly log-concave (even not necessarily independent) and have equal entropy $h_2(X)=h_2(Y)$.
\end{rmk}

\begin{thm}\label{thm:infty-ent-jump}
Let $X$ and $Y$ be independent real-valued log-concave random variables. Let $Z$ and $W$ be independent exponential random variables such that $h_\infty(Z)=h_\infty(X)$ and $h_\infty(W)=h_\infty(Y)$. We have
\begin{align}\label{eq:infty-ent-jump-a}
h_\infty(X+Y)\leq h_\infty(Z+W).
\end{align}
Consequently, for i.i.d. real-valued log-concave random variables $X$ and $Y$, we have
\begin{align}\label{eq:infty-ent-jump-b}
h_\infty(X+Y)\leq h_\infty(X)+1.
\end{align}
\end{thm}

The decreasing rearrangement operation tends to increase concentration on one side while reducing the additive overlap of translates in the convolution. Consequently, one expects $f^\downarrow\ast f^\downarrow$ to be less concentrated than $f\ast f$. This behavior is illustrated by the following result, which can be viewed as a reverse of Riesz's rearrangement inequality. The proof draws from our communication with James Melbourne.

\begin{lem}\label{lem:unimodal}
Let $f, g: \R\rightarrow \R_+$ be integrable unimodal functions. Then it holds that
$$
\|f\ast g\|_\infty\geq \|f^\downarrow\ast g^\downarrow\|_\infty.
$$
\end{lem}

\begin{proof}

For any given $z\ge 0$, the unimodality of $f$ and $g$ implies that there exist two intervals $A=[\alpha, \alpha+z]$ and $B=[\beta, \beta+z]$ such that 
$$
(f_{A})^\downarrow (y)=f^\downarrow (y),\quad (g_{B})^\downarrow (y)=g^\downarrow (y), \quad \text{for all}~y\in [0, z].
$$ 
(Here, we write $f_E$ and $g_E$ for the restrictions of $f$ and $g$ to $E\subseteq\R$, respectively). For $y\in A$, we define $\tilde{g}(y)=g_B(\gamma_z-y)$, where $\gamma_z=\alpha+\beta+z$, so that both $\tilde{g}$ and $f_A$ are supported on $A$.. Note that reflection and translation preserve the decreasing rearrangement of a function. Thus we have  
$$
(\tilde{g})^\downarrow(y)=(g_B)^\downarrow(y),\quad \text{for all}~y\in [0, z].
$$
Now we apply Lemma \ref{lem:rearrangement} to obtain
\begin{align*}
(f^\downarrow\ast g^\downarrow)(z) &=\int_0^zf^\downarrow(y)g^\downarrow(z-y)dy\\
&=\int_0^z(f_A)^\downarrow(y)(g_B)^\downarrow(z-y)dy\\
&=\int_0^z(f_A)^\downarrow(y)(\tilde{g})^\downarrow(z-y)dy\\
&\le \int_\alpha^{\alpha+z} f_A(y)\tilde{g}(y)dy\\
&=\int_\alpha^{\alpha+z}f_A(y)g_B(\gamma_z-y)dy\\
&= \int_\alpha^{\alpha+z} f(y)g(\gamma_z-y)dy\\
&\le(f\ast g)(\gamma_z).
\end{align*}
Then we can obtain the result by taking the supremum over $z\in\R$.
\end{proof}

\begin{rmk}
As shown in the proof, unimodality guarantees the existence of intervals $A$ and $B$ such that $(f_{A})^\downarrow=f^\downarrow$ and $(g_{B})^\downarrow=g^\downarrow$ on $[0, z]$. The result may fail for arbitrary functions that lack this interval structure. For example, define $f=\mathbf{1}_{[0, 2]}$ and $g=\mathbf{1}_{[0, 1]}+\mathbf{1}_{[3, 4]}$. Clearly, $f^\downarrow=g^\downarrow=\mathbf{1}_{[0, 2]}$. One can check that $\|f\ast g\|_\infty=1$ while $\|f^\downarrow\ast g^\downarrow\|_\infty=2$.
\end{rmk}

\begin{lem}\label{lem:exponentialization}
Let $X$ and $Y$ be independent monotone log-concave random variables. Let $Z$ and $W$ be independent exponential random variables such that $h_\infty(Z)=h_\infty(X)$ and $h_\infty(W)=h_\infty(Y)$. For all $0<p\leq\infty$, we have
\begin{align}\label{eq:major-exp}
h_p(X+Y)\leq h_p(Z+W).
\end{align}
\end{lem}

\begin{proof}
We will show the majorization of $X+Y$ over $X+W$ (i.e., the density of $X+Y$ majorizes that of $X+W$) and, in the same vein, the majorization of $X+W$ over $Z+W$. Hence, $Z+W$ is majorized by $X+Y$. This, together with inequality \eqref{eq:maj}, yields the statement \eqref{eq:major-exp}. 

Now, we show the majorization of $X+Y$ over $X+W$. Note that $X+Y$ and $X+W$ are log-concave. By Proposition \ref{prop:maj-suf}, it suffice to show that for any $[a, a+\delta]\subseteq [0, \infty)$, there exists $[b, b+\delta]\subseteq [0, \infty)$ such that
\begin{equation}\label{eq:maj-prob}
\P(X+W \in [a, a+\delta])\leq  \P(X+Y \in [b, b+\delta]).
\end{equation}
Let $f_X$ and $f_Y$ be the densities of $X$ and $Y$, respectively. We can perturb $f_X$ and $f_Y$ if necessary such that both of them have infinite support. We can also make shifts if necessary so that $\|f_X\|_\infty=f_X(0)$ and $\|f_Y\|_\infty=f_Y(0)$. Write $f_Y(x)=f_Y(0)e^{-V(x)}$ for $x\geq 0$, where $V(x)$ is an non-decreasing convex function with $V(0)=0$. Then $W$ has density $f_W(x)=f_Y(0)e^{-f_Y(0)x}$ for $x\geq0$. Consider the map $\phi =F_W^{-1}\circ F_Y$, where $F_Y$ and $F_W$ are the cumulative distribution functions of $Y$ and $W$, respectively. One can check that $W\overset{d}{=}\phi(Y)$ and $\phi(y)=-f_Y(0)^{-1}\log(1-F_Y(y))$. Then we can rewrite inequality \eqref{eq:maj-prob} as
$$
\P(X+\phi(Y) \in [a, a+\delta])\leq  \P(X+Y \in [b, b+\delta]).
$$
By conditioning on $Y$, it suffices to show for all $y\geq0$ that
\begin{align}\label{eq:major-pt}
\P(X\in [a, a+\delta]-\phi(y))\leq  \P(X\in [b, b+\delta]-y).
\end{align}
Next, we demonstrate that this inequality holds for $b=\phi^{-1}(a)$. 
Since $Y$ is log-concave, the tail probability $1-F_Y(y)$ is also log-concave. Hence, $\phi$ is a convex map. Furthermore, we have
\begin{align}\label{eq:deriv-T}
\phi'(y)=\frac{e^{-V(y)}}{1-F_Y(y)}.
\end{align}
Since $\phi$ is convex, $\phi'$ is nondecreasing, thus $\phi'(y) \geq \phi'(0)=1$.

\textit{Case 1}:  $y\geq \phi^{-1}(a+\delta)$. Since $X$ is non-negative, inequality \eqref{eq:major-pt} trivially holds.

\textit{Case 2}: $\phi^{-1}(a)<y< \phi^{-1}(a+\delta)$. By integrating $\phi'$ over $[\phi^{-1}(a),y]$,  $\phi' \geq 1$ implies that
$y-\phi^{-1}(a)\leq \phi(y)-a$, which is equivalent to $a+\delta-\phi(y)\leq \phi^{-1}(a)+\delta-y$. Since $X$ is non-negative, we obtain
$$
\P(X\in [0, a+\delta-\phi(y)]\leq  \P(X\in [0, \phi^{-1}(a)+\delta-y]),
$$
which is inequality \eqref{eq:major-pt}. 

\textit{Case 3}: $y\leq \phi^{-1}(a)$. Similar to Case 2, we can apply $\phi' \geq 1$  to obtain $a-\phi(y)\geq \phi^{-1}(a)-y\geq 0$ by integrating $\phi'$ over $[y, \phi^{-1}(a)]$.  This, together with the monotonicity of the density of $X$, yields inequality \eqref{eq:major-pt}.
\end{proof}

\begin{proof}[Proof of Theorem \ref{thm:infty-ent-jump}]
Let $f_X$ and $f_Y$ be the densities of log-concave random variables $X$ and $Y$, respectively. As shown in Proposition \ref{prop:rearrange propt-function}, log-concavity is preserved by decreasing rearrangement, thus $f_X^\downarrow$ and $f_Y^\downarrow$ are also log-concave. Since decreasing rearrangement preserves the $L^p$ moments, $f_X^\downarrow$ and $f_Y^\downarrow$ are also probability densities such that $\|f_X^\downarrow\|_\infty=\|f_X\|_\infty$ and $\|f_Y^\downarrow\|_\infty=\|f_Y\|_\infty$. Let $X^{\downarrow}$ and $Y^{\downarrow}$ be random variables with densities $f_X^{\downarrow}$ and $f_Y^{\downarrow}$, respectively. Then we have
	$$
	h_{\infty}(X^{\downarrow})=h_{\infty}(X)=h_{\infty}(Z),~~~~ h_{\infty}(Y^{\downarrow})=h_{\infty}(Y)=h_{\infty}(W).
	$$
	By Lemma  \ref{lem:unimodal}, we have
	$$
	h_{\infty}(X+Y)\leq h_{\infty}(X^{\downarrow}+Y^{\downarrow}).
	$$
	By Lemma \ref{lem:exponentialization}, we have
	$$
	h_{\infty}(X^{\downarrow}+Y^{\downarrow}) \leq h_{\infty}(Z+W).
	$$
	Then we can combine these inequalities to obtain inequality \eqref{eq:infty-ent-jump-a}. Inequality \eqref{eq:infty-ent-jump-b} follows from simple calculations.
\end{proof}

\begin{rmk}
We give a short proof of the $p=\infty$ case of inequality \eqref{eq:major-exp} via measure transportation (in fact the change of variable). (This, together with Lemma \ref{lem:unimodal}, yields a short proof of Theorem \ref{thm:infty-ent-jump}). Since the map $\phi$ transports $Y$ to $W$, it holds for any function $\varphi$ integrable with respect to $F_W$ that
$$
\int_{\R_+} \varphi(\phi(x))f_Y(x)dx=\int_{\R_+} \varphi(x)f_W(x)dx.
$$
For any $y\geq 0$, we apply this identity to $\varphi(x)=f_X(y-x)$ for $0\leq x\leq y$ to obtain
\begin{align*}
(f_X\ast f_W)(y) &=\int_0^y f_X(y-x)f_W(x)dx\\
&= \int_0^{\phi^{-1}(y)} f_X(y-\phi(x))f_Y(x)dx\\
&\leq \int_0^{\phi^{-1}(y)} f_X(\phi^{-1}(y)-x)f_Y(x)dx\\
&=(f_X\ast f_Y)(\phi^{-1}(y)),
\end{align*}
where the inequality follows from the observation that $y-\phi(x)\geq \phi^{-1}(y)-x$, which is implied by \eqref{eq:deriv-T}, and the assumption that $f_X$ is non-increasing. This yields 
$$
h_\infty(X+Y)\leq h_\infty(X+W)
$$
and similarly we can obtain
$$
h_\infty(X+W)\leq h_\infty(Z+W).
$$
These two inequalities together yield the $p=\infty$ case of inequality \eqref{eq:major-exp}. One can see that the monotonicity of $f_X$ and the expansion property of the transportation map $\phi$ play the key role.
\end{rmk}

\begin{rmk}
Let $(X, Y)$ be an \textit{origin symmetric} log-concave random vector in $\R^2$. (The marginals are not necessarily independent). A result of Ball (\cite{Bal88}, Theorem 5) imply that 
$$
e^{h_\infty(X+Y)}\leq e^{h_\infty(X)}+e^{h_\infty(Y)}.
$$
This inequality can be seen as the entropic analog of Busemann's convexity theorem of intersection bodies \cite{Bus49}, which also motivates the study of reverse EPIs in \cite{BNT16, Li18}.
\end{rmk}


\section{Analogs for Interger-valued Log-concave Random Variables}


We first recall the majorization of two sequences. Let $\{f(k)\}_{k=1}^\infty$ and $\{g(k)\}_{k=1}^\infty$ be two summable non-negative sequences such that $\sum_{k=1}^\infty f(k)= \sum_{k=1}^\infty g(k)$. We say that $\{f(k)\}_{k=1}^\infty$ majorizes $\{g(k)\}_{k=1}^\infty$ if the decreasing rearrangements $\{f^\downarrow(k)\}_{k=1}^\infty$ and $\{g^\downarrow(k)\}_{k=1}^\infty$ satisfy that $\sum_{k=1}^n f^\downarrow(k)\geq \sum_{k=1}^k g^\downarrow(k)$ for all $n\in\Z_+$. Karamata's inequality (or majorization inequality) says that if $\{f(k)\}_{k=1}^\infty$ majorizes $\{g(k)\}_{k=1}^\infty$ then it holds for any convex function $\varphi$ on $\R$ that
\begin{equation}\label{eq:kar}
\sum_{k=1}^\infty (\varphi\circ f)(k)\geq \sum_{k=1}^\infty (\varphi\circ g)(k).
\end{equation}

An integer-valued random variable $X$ is called \textit{log-concave} if its probability mass function $\{f(k)\}_{k\in \Z}$ is a log-concave sequence; that is, $f(k+1)^2\geq f(k)f(k+2)$ holds for all $k\in\Z$ and the support of this sequence is a contiguous interval of $\mathbb{Z}$. Similar to \eqref{eq:renyi}, the order $p$ R\'enyi entropy of $X$ is defined as
$$
H_p(X)=\frac{1}{1-p}\log \sum_{k\in \Z} f(k)^p.
$$
Particularly, we have 
$H_\infty(X) =-\log \|f\|_\infty$.

\begin{prop}\label{prop:2-conv-disct}
Let $X$ and $Y$ be i.i.d log-concave random variables on $\Z$. It holds that
\begin{equation}
H_2(X+Y)<H_2(X)+\log 2.
\end{equation}
Moreover, the inequality cannot be improved.
\end{prop}

\begin{proof}
Let $\tilde{X}$ and $\tilde{Y}$ be independent copies of $X$ and $Y$, respectively. We denote by $\{f(k)\}_{k\in\Z}$ the probability mass function of  $X-\tilde{X}$. We define the geometric sequence $\{g(k)\}_{k\in\Z_+}$ as $g(k)=f(0)\left(\frac{1-f(0)}{1+f(0)}\right)^k$. One can check that $\sum_{k=0}^\infty f(k)=\sum_{k=0}^\infty g(k)$. Note that $f(0)=g(0)$ and $\{g(k)\}_{k\in\Z_+}$ is a log-affine sequence. We have the majorization of $\{f(k)\}_{k\in\Z_+}$ over $\{g(k)\}_{k\in\Z_+}$. We apply Karamata's inequality \eqref{eq:kar} to obtain
\begin{align*}
\sum_{k\in\Z} f(k)^2 &=2\sum_{k\in\Z_+} f(k)^2-f(0)^2\geq 2\sum_{k\in\Z_+} g(k)^2-f(0)^2 \\
&=\frac{1+f(0)^2}{2}\cdot f(0)>\frac{f(0)}{2}.
\end{align*}
The above inequality, together with the symmetry of $X-\tilde{X}$, implies that
\begin{align}\label{eq:3rd-b}
H_\infty((X-\tilde{X})+(Y-\tilde{Y})) &=-\log (f\ast f)(0) =-\log \sum_{k\in\Z} f(k)^2\notag\\
&<-\log f(0)+\log 2=H_\infty(X-\tilde{X})+\log 2.
\end{align}
Observe that for any i.i.d. discrete log-concave random variables $U$ and $V$, it holds that $H_2(U)=H_\infty(U-V)$. We have
\begin{align*}
H_2(X+Y) &=H_\infty((X+Y)-(\tilde{X}+\tilde{Y}))\\
&=H_\infty((X-\tilde{X})+(Y-\tilde{Y}))\\
&< H_\infty(X-\tilde{X})+\log 2\\
&= H_2(X)+\log 2.
\end{align*}
The inequality follows from \eqref{eq:3rd-b}. The moreover part is addressed later in Remark 4.4.
\end{proof}

\begin{rmk}
Similar to the continuous case, the identity $H_2(X+Y)=H_2(X-Y)$ holds for any $\Z$-valued i.i.d log-concave random variables $X$ and $Y$. Hence, Proposition \ref{prop:2-conv-disct} is the same as the $p=2$ case of Theorem 2.11 of Melbourne and Palafox-Castillo \cite{MPC23}. 
\end{rmk}

\begin{thm}\label{thm:infty-conv-disct}
Let $X$ and $Y$ be i.i.d. log-concave random variables on $\Z$. Assume that their probability mass function $f$ is monotone. It holds that
\begin{equation*}\label{eq:infty-conv-disct}
\|f\ast f\|_\infty> e^{-1}\|f\|_\infty.
\end{equation*}
Moreover, the inequality can not be improved. In other words, we have
$$
H_\infty(X+Y)<H_\infty(X)+1.
$$
\end{thm}

\begin{proof}
We assume that $f$ is non-increasing and  supported on $\Z_+$. Write $f(k)=e^{-V(k)}$, where $V$ is non-decreasing and convex. Hence, $\|f\|_\infty=e^{-V(0)}$. We apply the convexity of $V$ to obtain
\begin{align*}
(f\ast f)(k) 
&=\sum_{i=0}^ke^{-[V(i)+V(k-i)]}\geq \sum_{i=0}^ke^{-[V(0)+V(k)]}=(k+1)e^{-V(k)}\cdot \|f\|_\infty.
\end{align*}
Write $\phi(k):=(k+1)e^{-V(k)}=e^{\log(k+1)-V(k)}$. It suffices to show that
\begin{equation}\label{eq:max-disct}
\sup_{k\in\Z_+}\phi(k)>1/e.
\end{equation}
Since $\phi$ is log-concave, the supremum in \eqref{eq:max-disct} can be achieved, say at $k=k^*$. Consider the following scenarios.

\textit{Case 1}: $k^*=0$. Then it is necessary to have $\phi(0)\geq \phi(1)$, which can be rewritten as
$$
V(1)-V(0)\geq \log 2.
$$
Since $V$ is convex, it holds that
$$
V(k)\geq V(0)+k(V(1)-V(0))\geq V(0)+k\log 2.
$$
As a consequence, we obtain
$$
1=\sum_{k=0}^\infty f(k)=e^{-V(0)}\sum_{k=0}^\infty e^{-(V(k)-V(0))}\leq \phi(0)\sum_{k=0}^\infty 2^{-k}=2\phi(0). 
$$
Hence, it is necessary to have $\phi(0)\geq 1/2$ and inequality \eqref{eq:max-disct}  holds.

\textit{Case 2}: $k^*\geq 1$. Then it is necessary to have $\phi(k^*-1)\leq \phi(k^*)$ and $ \phi(k^*+1) \leq \phi(k^*)$, which can be rewritten as
\begin{align*}
V(k^*)-V(k^*-1)&\leq \log(1+1/k^*)\\
V(k^*+1)-V(k^*)&\geq \log(1+1/(k^*+1)).
\end{align*}
Owing to the convexity of $V$, we obtain for $k\leq k^*$ that
\begin{align}\label{eq:est-V-a}
V(k)&\geq V(k^*)-(k^*-k)[V(k^*)-V(k^*-1)]\notag\\
&\geq V(k^*)+(k-k^*)\log(1+1/k^*)
\end{align}
and for $k\geq k^*$ that
\begin{align}\label{eq:est-V-b}
V(k)&\geq V(k^*)+(k-k^*)[V(k^*+1)-V(k^*)]\notag\\
&\geq V(k^*)+(k-k^*)\log(1+1/(k^*+1)).
\end{align}
Apply the estimates \eqref{eq:est-V-a} and \eqref{eq:est-V-b} to obtain
\begin{align*}
1=\sum_{k=0}^\infty f(k)&\leq e^{-V(k^*)}\left(\sum_{k=0}^{k^*}\left(1+\frac{1}{k^*}\right)^{k^*-k}+\sum_{k=k^*+1}^\infty\left(1+\frac{1}{k^*+1}\right)^{-(k-k^*)}\right)\\
&=e^{-V(k^*)}\left(1+k^*\left(1+\frac{1}{k^*}\right)^{k^*+1}\right)\\
&=\phi(k^*)\left(\frac{1}{k^*+1}+\left(1+\frac{1}{k^*}\right)^{k^*}\right)\\
&<e\phi(k^*).
\end{align*}
In the last inequality, we use the fact that $(x+1)^{-1}+(1+1/x)^x\nearrow e$ as $x\rightarrow\infty$. Hence, it is necessary to have $\phi(k^*)>1/e$, i.e., inequality \eqref{eq:max-disct} holds.
\end{proof}

\begin{rmk}
Proposition \ref{prop:2-conv-disct} and Theorem \ref{thm:infty-conv-disct} are tight for geometric distributions $f(k)=(1-\lambda)\lambda^k$ for $k\in\Z_+$. One can check that $(f\ast f)(k)=(k+1)(1-\lambda)^2\lambda^k$ and that
$$
e^{H_2(f\ast f)-H_2(f)}=\frac{\sum_{k=0}^\infty f(k)^2}{\sum_{k=0}^\infty (f\ast f)(k)^2}=\frac{(1+\lambda)^2}{1+\lambda^2}\rightarrow 2,~~\text{as}~\lambda\rightarrow 1.
$$
It is clear that $\|f\|_\infty=1-\lambda$. Set $\lambda=e^{-1/(k+1)}$. As $k\rightarrow\infty$, we have
$$
\frac{\|f\ast f\|_\infty}{\|f\|_\infty}\geq (k+1)(1-\lambda)\lambda^k=(k+1)(1-e^{-1/(k+1)})e^{-k/(k+1)}\rightarrow \frac{1}{e}.
$$
If the probability mass function of $X$ is symmetric about an integer or a half-integer, we can apply the argument in Proposition \ref{prop:2-conv-disct} to establish $H_\infty(X+Y)<H_\infty(X)+1/2$.
\end{rmk}

{\bf Acknowledgement.} J. L. would like to thank Mokshay Madiman and James Melbourne for valuable discussions. The work is supported by the National Natural Science Foundation of China (NSFC) grant 62201175.


\bibliographystyle{plain}

\end{document}